### v1

# Quantum Lamb model

Abstract
H. Lamb considered the classical dynamics of a vibrating particle embedded in an elastic medium before the development of quantum theory. Lamb was interested in how the back-action of the elastic waves generated can damp the vibrations of the particle. We propose a quantum version of Lamb’s model. We show that this model is exactly solvable by using a multimode Bogoliubov transformation. We show that the exact system ground state is a multimode squeezed vacuum state, and we obtain the exact Bogoliubov frequencies by numerically solving a nonlinear integral equation. A closed-form expression for the damping rate of the particle is obtained, and we find that it agrees with the result obtained by perturbation theory for coupling strength below a critical value. The model is found to break down for coupling strength above the critical value where the lowest Bogoliubov frequency vanishes. We show that the addition of an anharmonic elastic term is sufficient to stabilize the system in this strong coupling regime.
preprint: arXiv:
I Introduction
Advances in the fabrication and characterization of simple mechanical systems in the nanoscopic or mesoscopic regime have driven experimental and theoretical investigations of some of the foundational principles of quantum mechanics. Prominent classes of such systems include vibrating beams and mirrored surfaces that interact with laser light through its radiation pressure (optomechanics) [1], mechanical resonators coupled to electronic devices (nanoelectromechanics) [2], and interacting mechanical resonators (quantum acoustodynamics) [3, 4, 5]. In addition to providing a path to explore quantum science and the limits of precision measurement, such systems might be used to fashion new quantum sensors and devices for manipulating quantum information, ushering in new quantum technologies [6, 7].
We consider a mechanical system whose first study predates the development of quantum mechanics. In 1900, Lamb [8] considered the dynamics of a vibrating particle embedded in an elastic medium. The back-action of the elastic waves generated by the vibrations of the particle work to damp those vibrations. In this work, we study a quantum version of Lamb’s model and focus on the dynamics of the vibrational decay. Figure 1 shows a schematic of the model consisting of a vibrating bead coupled to a light, long string under tension.
The proposed quantum Lamb model is exactly solvable with the use of a multimode Bogoliubov transformation. We find explicit expressions for the coefficients that diagonalizes the Hamiltonian. Using the symplectic properties of the transformation, we confirm that our results satisfy the necessary identities. We then derive a nonlinear equation whose solution yields the Bogoliubov frequencies, and we use it to numerically calculate the symplectic spectrum of the model.
We show that the ground state of the quantum Lamb model is a nonclassical state–a multimode squeezed vacuum state– and we relate this ground state to the uncoupled states (transverse phonons of the string and vibrons of the bead) of the system. We then study the dynamics of the vibrational decay of the bead. We obtain an explicit expression for the decay rate, and we calculate the spectral distribution of single bogoliubon emission, a product of the decay. We show that for weak coupling, the decay rate calculated agrees with both the classical result and the golden rule. For sufficiently strong coupling, the model becomes unstable where the lowest Bogoliubov frequency vanishes; however, the model can be stabilized in the strong coupling regime by including the anharmonic contribution of the bead’s elastic energy to the Hamiltonian.
$m$$u_{0}(t)$$x$
Figure 1: Schematic of a possible configuration for the classical Lamb model. Bead of mass $m$ at $x=0$ is constrained to move in the vertical direction and feels a restoring spring force and a contact force exerted by long, light string under tension $\tau$. The vibrating bead creates transverse acoustic waves on the string ($\ell\gg c/\omega_{0}$).
II Hamiltonian
We take for the Hamiltonian
$H=\sum_{\alpha=0}^{N}\omega_{\alpha}a_{\alpha}^{\dagger}a_{\alpha}-\left(a_{0}+a_{0}^{\dagger}\right)\sum_{n=1}^{N}\gamma_{n}\left(a_{n}+a_{n}^{\dagger}\right)$ (1)
where $a_{n}^{\dagger}$ ($a_{n}$) creates (annihilates) a transverse acoustic phonon on the string, and $a_{0}^{\dagger}$ ($a_{0}$) creates (annihilates) a vibron on the bead. (We use index notation where (greek) $\alpha=0,1,2,\dots$, while (roman) $n=1,2,\dots,N$, and work with natural units where $\hbar=1$.)
$\omega_{0}$ is the uncoupled vibrational frequency of the bead, while $\omega_{n}$ are the vibrational frequencies of the string (tension $\tau$, length $\ell$, transverse speed of sound $c$), subject to homogeneous boundary conditions $u_{s}(x=0,t)=0$ and $\partial_{x}u_{s}|_{x=\ell}=0$.
The coupling parameters $\gamma_{n}$ can be expressed in terms of physical quantities of the model, with $\gamma_{n}=\sqrt{\nu\omega_{n}{c}/{\ell\omega_{0}}}$ and $\nu\equiv\frac{\tau}{2mc}$.
III Bogoliubov transformation
We diagonalize the Hamiltonian in Eq. 1 with the use of a multimode Bogoliubov transformation. We look for a linear transformation (and its inverse) with the following form:
$b_{\alpha}$ $=$ $\sum_{\beta}\left(M_{\alpha\beta}a_{\beta}+N_{\alpha\beta}a_{\beta}^{\dagger}\right)$ (2)
$a_{\alpha}$ $=$ $\sum_{\beta}\left(U_{\alpha\beta}b_{\beta}+V_{\alpha\beta}b_{\beta}^{\dagger}\right)$ (3)
where $b_{\alpha}^{\dagger}$ ($b_{\alpha}$) creates (destroys) a Bogoliubov excitation (bogoliubon) and M, N, U, and V are $(N+1)$-dimensional square matrices whose elements are the coefficients of the transformation.
We require that the transformation preserve the boson commutation rules $\left[b_{\alpha},b^{\dagger}_{\beta}\right]=\delta_{\alpha\beta}$. As a result [9, 10], the coefficients can be grouped to form a $2(N+1)$-dimensional symplectic matrix $\textsf{T}\in\text{Sp}(2(N+1),\mathbb{R})$:
$\textsf{T}\equiv\begin{pmatrix}\textsf{M}&\textsf{N}\\ \textsf{N}&\textsf{M}\end{pmatrix}.$ (4)
T then satisfies the symplectic condition $\textsf{T}\textsf{J}\textsf{T}^{T}=\textsf{J}$ where the symplectic form J can be represented as
J( $\textsf{J}\equiv\begin{pmatrix}\ \ 0\kern-3.99994pt\text{0}&1\kern-2.5pt\text{l}\\ -1\kern-2.5pt\text{l}&0\kern-3.99994pt\text{0}\end{pmatrix}.$ (5)
A number of useful coefficient identities follow [11] from the symplectic structure on this Fock space; for example, the inverse of the transformation matrix T can be obtained simply from the symplectic condition (together with $\textsf{J}^{2}=-1\kern-2.5pt\text{l}$):
$\textsf{T}^{-1}$ $=$ $\textsf{T}^{T}\textsf{J}$ J (6)
$=$ $\begin{pmatrix}\textsf{M}^{T}&-\textsf{N}^{T}\\ -\textsf{N}^{T}&\textsf{M}^{T}\end{pmatrix}.$ (7)
Hence, we conclude that the coefficients of the inverse transformation satisfy $\textsf{U}=\textsf{M}^{T}$ and $\textsf{V}=-\textsf{N}^{T}$. We summarize the explicit form for the transformation in Table 1. The detailed calculation of the coefficients is outlined in the supplemental material [11].
Table 1: Coefficients for the Bogoliubov transformation $M_{\alpha\beta}$ and $N_{\alpha\beta}$ ($\alpha,\beta=0,1,\dots,N$ and $k,q=1,2,\dots,N$). Coefficients for the inverse transformation can be obtained from the transpose relations $U_{\alpha\beta}=M_{\beta\alpha}$ and $V_{\alpha\beta}=-N_{\beta\alpha}$ (see supplemental material [11].
$(\alpha 0)$ $(\alpha k)$
$M$ $\frac{\Omega_{\alpha}+\omega_{0}}{\sqrt{4\omega_{0}\Omega_{\alpha}}}\frac{1}{\sqrt{1+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{\left(\Omega_{\alpha}^{2}-\omega_{q}^{2}\right)^{2}}}}$ $-\frac{2\omega_{0}\gamma_{k}}{(\Omega_{\alpha}-\omega_{k})}\frac{1}{\sqrt{4\omega_{0}\Omega_{\alpha}}}\frac{1}{\sqrt{1+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{\left(\Omega_{\alpha}^{2}-\omega_{q}^{2}\right)^{2}}}}$
$N$ $\frac{\Omega_{\alpha}-\omega_{0}}{\sqrt{4\omega_{0}\Omega_{\alpha}}}\frac{1}{\sqrt{1+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{\left(\Omega_{\alpha}^{2}-\omega_{q}^{2}\right)^{2}}}}$ $-\frac{2\omega_{0}\gamma_{k}}{(\Omega_{\alpha}+\omega_{k})}\frac{1}{\sqrt{4\omega_{0}\Omega_{\alpha}}}\frac{1}{\sqrt{1+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{\left(\Omega_{\alpha}^{2}-\omega_{q}^{2}\right)^{2}}}}$
The following Hamiltonian results
$H=\sum_{\alpha}\Omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha}$ (8)
where the Bogoliubov frequencies $\{\Omega_{\alpha}\}$ satisfy the following nonlinear equation:
$\Omega_{\alpha}^{2}=\omega_{0}^{2}+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{\Omega_{\alpha}^{2}-\omega_{q}^{2}}.$ (9)
We note that the Hamiltonian is no longer positive-definite when the lowest Bogoliubov frequency vanishes. Thus there is a constraint on the model; namely,
$\frac{4}{\omega_{0}}\sum_{q}{\frac{\gamma_{q}^{2}}{\omega_{q}}}<1.$ (10)
We define the coupling strength $g\equiv\frac{4}{\omega_{0}}\sum_{q}{\frac{\gamma_{q}^{2}}{\omega_{q}}}$ and conclude that there is a critical coupling strength $g_{c}=1$ above which the model is ill-defined. Figure 2 shows a plot of the Bogoliubov frequencies $\Omega_{\alpha}/\omega_{0}$ obtained numerically as a function of the coupling strength $g$ for the lowest 10 modes.
Figure 2: Bogoliubov frequencies $\Omega_{\alpha}/\omega_{0}$ versus coupling strength $g$ for the lowest 10 frequencies. $\Omega_{1}$ vanishes at the critical point $g=1$, signaling a quantum phase transition. In this plot, we have set $\Omega_{0}/\omega_{0}=1$ for $g=0^{+}$, and the string spectrum is taken to be $\omega_{k}/\omega_{0}=\epsilon+k\delta$ with $\epsilon=0.1$, $\delta=0.25$, $k=1,2,\dots,N$ with $N=9$.
IV Multimode squeezed vacuum
Eigenstates of the system can be labeled by the set of Bogoliubov excitation numbers for the $N+1$ modes |$|\{n_{\alpha}\}\rangle$ with corresponding energies $E(\{n_{\alpha}\})=\sum_{\alpha}n_{\alpha}\Omega_{\alpha}+\frac{1}{2}\sum_{\alpha}(\Omega_{\alpha}-\omega_{\alpha})$. The ground state of the coupled system |$|\{0\}\rangle$ can be constructed from the uncoupled ground state $|\rangle_{0}$ with the squeeze operator $S({\xi})=\exp\left(-\frac{1}{2}\sum_{\alpha\beta}\xi_{\alpha\beta}a^{\dagger}_{\alpha}a^{\dagger}_{\beta}\right)$:
|$|\{0\}\rangle={\cal N}\ S({\xi})|\rangle_{0}.$ (11)
$\xi$ is the (matrix) squeeze parameter and ${\cal N}$ is a normalization factor. (Using an identity due to Schwinger [12], the normalization constant is ${\cal N}=\frac{1}{\sqrt{\det\textsf{M}}}$.)
We verify this by operating on Eq. 11 with $b_{\alpha}$ and using the identity $S(-{\xi})a_{\alpha}S({\xi})=a_{\alpha}-\sum_{\beta}\xi_{\alpha\beta}a_{\beta}^{\dagger}$. We find that Eq. 11 is satisfied, provided the squeeze matrix has the value
$\xi=\textsf{M}^{-1}\ \textsf{N}.$ (12)
Hence, the ground state of the model is a multimode squeezed vacuum state [13, 9].
The average number of uncoupled excitations (phonons and vibrons) in the mode $\alpha$ contained in the coupled ground state |$|\{0\}\rangle$ can be expressed in terms of Bogoliubov coefficients, with
$n_{\alpha}=\sum_{\beta}N_{\beta\alpha}^{2}.$ (13)
An example is given in Fig. 3 for a coupling strength of $g=0.7$. There is a small fraction of a vibron contributed by the bead ($\alpha=0$), with an equal total amount of phonons on the string approximately uniformly distributed across the modes at this coupling strength.
Figure 3: Distribution of the average number of uncoupled excitations (phonons and vibrons) in the coupled ground state $n_{\alpha}=\langle\{0\}|a^{\dagger}_{\alpha}a_{\alpha}|\{0\}\rangle$ for parameter values $N=15$, $g=0.7$.
V Vibron decay
We now consider the dynamics of the vibrational decay of the bead. We start in the ground state |$|\{0\}\rangle$, displace the bead by $\delta$ to create the initial state $|\Psi(0)\rangle=\exp(-ip_{0}\delta)|\{0\}\rangle$, and compute the expectation of the bead’s position at time $t$:
$\langle u_{0}(t)\rangle=\langle\Psi(t)|u_{0}|\Psi(t)\rangle.$ (14)
For $\delta\ll\sqrt{\hbar/m\omega_{0}}$,
$\langle u_{0}(t)\rangle$ $=$ $\delta\ {\rm Re}\ \sum_{\alpha}\left(U_{0\alpha}^{2}-V_{0\alpha}^{2}\right)\exp(-i\Omega_{\alpha}t)$ (15)
$=$ $\delta\ {\rm Re}\ \sum_{\alpha}{\frac{\exp(-i\Omega_{\alpha}t)}{1+4\omega_{0}\sum_{n}{\frac{\gamma_{n}^{2}\omega_{n}}{\left(\Omega_{\alpha}^{2}-\omega_{n}^{2}\right)^{2}}}}}.$ (16)
We identify the factor $\left(U_{0\alpha}^{2}-V_{0\alpha}^{2}\right)$ in the summand of Eq. 15 as the spectral density of the decay:
$\rho(\Omega_{\alpha})=U_{0\alpha}^{2}-V_{0\alpha}^{2}.$ (17)
This spectral density satisfies a sum rule [11] $\sum_{\alpha}\rho(\Omega_{\alpha})=1$, and the width of this spectral density is the decay rate of the bead displacement [14] (see Fig. 4).
Figure 4: Spectral distribution function $\rho(\Omega_{\alpha})$ versus $\Omega_{\alpha}/\omega_{0}$ for $g=0.4$ and $0.6$. It satisfies the sum rule $\sum_{\alpha}\rho(\Omega_{\alpha})=1$ and its width gives the decay rate $\Gamma$ of $\langle u_{0}(t)\rangle$.
Using contour integration in the complex plane, the sum can be evaluated [11] and the decay rate $\Gamma$ can be obtained:
$\Gamma=\frac{\omega_{r}}{\sqrt{2}}\left(\sqrt{1+\frac{\Gamma_{r}^{4}}{\omega_{r}^{4}}}-1\right)^{\frac{1}{2}}$ (18)
where $\omega_{r}\approx\omega_{0}$ and $\Gamma_{r}\approx\left(\frac{\tau\omega_{0}}{2mc}\right)^{\frac{1}{2}}$. For the case of light damping where $\Gamma_{r}\ll\omega_{r}$, Eq. 18 gives $\Gamma\approx\frac{\tau}{2mc}=\nu$, in agreement with the classical result [11]. For $\Gamma_{r}\gg\omega_{r}$, $\Gamma$ behaves as $\Gamma\to\sqrt{\frac{\nu\omega_{0}}{2}}$.
We now turn to the radiation spectrum from the vibrating bead. The decay of the vibrating bead is accompanied by the emission of bogoliubons. The probability of the emission of a single bogoliubon of frequency $\Omega_{\alpha}$ can be expressed in terms of Bogoliubov coefficients:
$P_{1}(\Omega_{\alpha})=|\langle\{0\}|b_{\alpha}a_{0}^{\dagger}|\rangle_{0}|^{2}=M^{-2}_{0\alpha}\ (\det{M})^{-1}.$ (19)
A plot of the spectral probability distribution for single bogoliubon emission is given in Fig. 5.
Figure 5: Spectral probability distribution for single bogoliubon emission $P_{1}(\Omega_{\alpha})$ from the decay of a vibron $|1;\{0\}\rangle$. Parameter values are $g=0.7$ and $N=15$. The spectrum of single bogoliubons comprise 90.7% of the total emission.
VI Strong coupling
The vanishing Bogoliubov frequency at $g=1$ signals an unstable ground state for the model. In an attempt to stabilize the model for coupling strengths $g>1$, we add to Eq. 1 a quartic elastic energy term for the bead displacement $H^{\prime}=\frac{\lambda}{4}u_{0}^{4}$.
We find that it is energetically favorable for the classical equilibrium position of the bead to be displaced for $g>1$. Parity is spontaneously broken for $g>1$. We then consider small displacements about the displaced equilibrium position. The bead frequency $\omega_{0}$ is then renormalized, with $\bar{\omega}_{0}^{2}=\omega_{0}^{2}(3g^{2}-2)$. The coupling parameters $\gamma_{n}$ are also renormalized, as they depend on the bead frequency with $\bar{\gamma}_{n}\propto\bar{\omega}_{0}^{-\frac{1}{2}}$ (see Ref. [11]).
The spectrum of the Hamiltonian with renormalized parameters is then calculated, and we find that the renormalized Hamiltonian is positive-definite in the strong coupling regime $g>1$. Figure 6 shows the spectrum for the anharmonic model for $0\leq g\leq 2$.
Figure 6: Spectrum of Bogoliubov frequencies $\Omega_{\alpha}/\omega_{0}$ versus coupling strength $g$ for $0\leq g\leq 2$. Lowest 10 frequencies are displayed. Hamiltonian contains a quartic contribution to the elastic energy of the bead that stabilizes the model for $g>1$.
VII Discussion
We analyzed the dynamics of a vibrating particle coupled to an environment by extending the Lamb model to the quantum regime. The model has the virtue that it is exactly solvable using a multimode Bogoliubov transformation. We found that the coupled ground state of the model is a multimode squeezed vacuum state, and we obtained the bogoliubon spectrum. We then obtained an explicit expression for the vibrational decay rate of the bead and found that it recovered the classical damping rate in the light damping regime.
We examined the acoustic radiation spectrum emitted by the vibrating particle. We obtained an expression for the probability of single bogoliubon emission and observed that it has a nearly symmetric lineshape about a slightly red-shifted frequency.
The linear model is unstable at a critical coupling strength $g_{c}=1$; however, we find that the stability can be restored with the inclusion of a quartic elastic energy term in the Hamiltonian. The parity symmetry of the model is spontaneously broken at strong coupling, and the bead frequency for oscillations about the symmetry-broken equilibrium is stiffened with increasing coupling strength. A positive-definite Hamiltonian results for $g\neq 1$ by including the elastic anharmonicity into the linear Lamb model. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) and grant NASA 80NSSC19M0143.